OpenStudy (anonymous):

The ordered pairs (1, 1), (2, 8), (3, 27), (4, 64), and (5, 125) represent a function. What is a rule that represents this function?

OpenStudy (anonymous):

f(x)=1 (x=1)/ 8 (x=2)/ 27 (x=3)/ 64 (x=4) / 125 (x=5)

OpenStudy (anonymous):

or f(x) = x^3